OpenStudy (anonymous):

Solve Z = 9x + 9y for y.

OpenStudy (anonymous):

I know it's either y=z-x divided by 9 or y=z-9x divided by 9

OpenStudy (shane_b):

This is just shifting things around:\[Z=9x+9y\]\[Z-9x=9y\]\[\frac{Z-9x}{9}=y\]

OpenStudy (anonymous):

K I was just wondering we would not divide the -9x with 9 right?

OpenStudy (shane_b):

You could do that...but you'd have to do it with Z as well:\[\frac{Z}{9}-x=y\]

OpenStudy (anonymous):

Oh okay. That's not in my choices so yeah it's the one you wrote first. :) Tyvm.

OpenStudy (shane_b):

no worries..good luck